the mass of Jupiter is 1.901x1027 kg and its radius is 7.144x107m. calculate the value of gravitational acceleration, g, on the surface of jupiter


1
Expert's answer
2021-10-24T18:27:56-0400

We can find the gravitational acceleration from the formula:


"g_J=\dfrac{GM_J}{R_J^2},""g_J=\dfrac{6.67\times10^{-11}\ \dfrac{N\times m^2}{kg^2}\times1.901\times10^{27}\ kg}{(7.144\times10^7\ m)^2}=24.8\ \dfrac{m}{s^2}."